14. Henry Sold 8 more plants than Diane. Together they sold 56 plants. write and solve an equation to answer these questions: How Many plants did Diane Sell? How many plants did Henry Sell?
Answer by richwmiller(17219)

h=d+8,
h+d=56
d+8+d=56
2d=48
d=24
h=32